OpenStudy (anonymous):

2e^5x=18 solve please

OpenStudy (anonymous):

Isolate x so just divide everything by 2e^5 and you will get your answer \[\frac{2e^5x}{2e^5}= \frac{18}{2e^5}\] \[x=\frac{18}{2e^5}\]

OpenStudy (anonymous):

Thank you!! Could you please answer another. 5^x+1=3

OpenStudy (anonymous):

Ok so again we need to try to isolate the x So subtract one from both sides

OpenStudy (anonymous):

what do ya get?

OpenStudy (anonymous):

I am isolating them.. But I am not getting the correct answer whe I check them

OpenStudy (anonymous):

One step at a time. First subtract 1 from both sides. What do you get?

OpenStudy (anonymous):

-.4 is what I get for x

OpenStudy (anonymous):

And this not = 3. Unless I was making an error somewhere.

OpenStudy (anonymous):

\(5^x+1=3\) \(5^x+1-1=3-1\) \(5^x=2\)

OpenStudy (anonymous):

Do you have solutions?

OpenStudy (anonymous):

Have you learned about the ln function?

OpenStudy (anonymous):

Yes I have learned about the ln function and log. but my proffesor was very quick when he taught it to us.. Plus he is russian. The +1 is also sqaured.

OpenStudy (anonymous):

makes no diff if 1 is squared cuz 1*1*1*1*1*....*1=1 right?

OpenStudy (anonymous):

but I dont understand how you can subtract a one over to the other side when you have 5^x+1. What do you do with the 5?

OpenStudy (anonymous):

Hard to explain over this lol. Im not being very clear

OpenStudy (anonymous):

in my first question. I took 18/2e^5 and got .0606 and that did not come out right

OpenStudy (anonymous):

well can you draw the question to make sure i have the right equation?

OpenStudy (anonymous):

I dont know how to make it but i can only do my best. its 5^x+1=3. The 0ne is included with the x pwer.

OpenStudy (anonymous):

I tried the logginng by two method, but I dont think that works

OpenStudy (anonymous):

Fine got the second equation but what abt the original equation. The first one I solved. It was incorrect? Also there is a drawing function right under the textbox

OpenStudy (anonymous):

 This was what I got. I punched it in and got .06.. Which doesnt check out right. Original equation is